OpenStudy (anonymous):

A geologist at point A measures the distance and angle of elevation to two columns of the Glen Rose formation in Texas. If the distance AS = 790ft and the distance AN = 883ft, compute the difference in the heights of the two columns to the nearest foot. [A] 1 ft [B] 970 ft [C] 260 ft [D] 118 ft

OpenStudy (anonymous):

I am so confused... Idk what to do...

OpenStudy (anonymous):

Is there a picture that goes with this?

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

sin(49) = length of SB/length of AS sin(54) = length of NC/length of AN

OpenStudy (anonymous):

oh ok..

OpenStudy (anonymous):

How would I find the length of SB?

Vocaloid (vocaloid):

what is the length of AS?

OpenStudy (anonymous):

790

Vocaloid (vocaloid):

great! so you can use my equation to find the length of SB sin(49) = length of SB/length of AS

Vocaloid (vocaloid):

put in the length of AS and solve for SB

OpenStudy (anonymous):

Ok, let me try.

Vocaloid (vocaloid):

make sure your calculator is in degree mode

OpenStudy (anonymous):

Ok. Would I divide AS by sin(49?

OpenStudy (anonymous):

How would that work?

Vocaloid (vocaloid):

not quite: here's a hint: a = b/c b = c*a

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

596.2?

Vocaloid (vocaloid):

great! now do the same thing for the second equation: sin(54) = length of NC/length of AN

OpenStudy (anonymous):

Sorry I was eating dinner

OpenStudy (anonymous):

714.3

OpenStudy (anonymous):

So the answer (final) would be D, correct?

Vocaloid (vocaloid):

yup! good work!

OpenStudy (anonymous):

thank you!